OpenStudy (konradzuse):

The population (in millions) of a country in 2007 and the expected continuous annual rate of change k of the population are given. † Country 2007 Population k Latvia 2.3 − 0.006 (a) Find the exponential growth model P = Cekt for the population by letting t = 0 correspond to 2000. (b) Use the model to predict the population of the country in 2011. (Round your answer to two decimal places.) million

OpenStudy (konradzuse):

P = Ce^kt = Ce^(-0.006) p(7) = 2.3 =Ce^-0.006(7) C = 2.40? for t = 11 P = 2.40e^(-0.006(11)) = 2.25 million?

OpenStudy (konradzuse):

cool works.